OpenStudy (anonymous):

I'm soo lost!!! HELP please Q: The board members of a provincial organization receive a car allowance for travel meetings. Here are the distances the board logged last year (in kilometres) 44, 18, 125, 80, 63, 42, 35, 68, 52, 75, 260, 96, 110, 72, 51. Determine the standard deviation, variance, interquartile range, and semi interquartile range.

OpenStudy (amilapsn):

do you know the way you should find the standard deviation?

OpenStudy (anonymous):

no i have no idea

OpenStudy (anonymous):

i don't know what x-mu is support to substitute for

OpenStudy (anonymous):

I'll write down the formulas for you in case you don't know them , and if you have further questions you can ask them standard deviation = \[\sqrt{variance}\] variance = \[\sqrt{(Sum ^{2})/number of terms)-(Sum/number of terms)^2}\]

OpenStudy (anonymous):

To find the interquartile range, you need to get both the upper quartile, and the lower quartile \[3n/4\] <---- upper quartile

OpenStudy (anonymous):

n/4 <--- lower quartile where n = number of terms

OpenStudy (anonymous):

after you get an answer for upper quartile, say for example the answer is 4, then your upper quartile final answer would be the 4th number in the given set of numbers

OpenStudy (anonymous):

and you do the same for lower quartile, after which to get the interquartile range, you subtract : upper quartile - lower quartile

OpenStudy (anonymous):

now how about you try finding the answers, and if you still get them wrong then I can guide you from there

OpenStudy (anonymous):

thank you very much, i can't seem to get the standard deviation down though. my book is telling me to do x-mu to get the deviation and to get the variation it says to do (x-mu)^2 and i don't know what to substitute mu for. So the first one would be 44-mu<-- i have no clue how to get this number. For example, one of the numbers on the example question says the standard deviation of 183 is 18

OpenStudy (anonymous):

well um I personally have never seen that equation before :P i dont know what m and u are either :/

OpenStudy (anonymous):

i circled it, its called sigma

OpenStudy (anonymous):

the picture is mirrored :/

OpenStudy (anonymous):

this is what mu means ---> μ the equation says its x-μ

OpenStudy (anonymous):

i figured it out, the mean is μ and x is the data number

OpenStudy (anonymous):

oh okay, so can you solve the question now?

OpenStudy (anonymous):

yes, thanks for all your help

OpenStudy (anonymous):

you're welcome